OpenStudy (anonymous):

simplify 9a²-b² ------- ab _________ 3a+b ----- a 3a+b a

OpenStudy (anonymous):

disreguard the last 3a+b a it is typo

OpenStudy (anonymous):

A/B/C/D=AD/BC

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

Consider the first line of your problem A, second B and so on. You can put the first, last on top, two middle bottom

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

well, well, what have we here>>>

OpenStudy (anonymous):

do u know the answer?

OpenStudy (anonymous):

cld u type out the question again ?

OpenStudy (anonymous):

simplify 9a²-b² ------- ab _________ 3a+b ----- a

OpenStudy (anonymous):

its a difference of square

OpenStudy (anonymous):

o.k, so does that mean divide by ab?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(3a+b)(3a-b)

OpenStudy (anonymous):

(3a-b)/b

OpenStudy (anonymous):

oh, i see its a complex fraction

OpenStudy (anonymous):

so ur a is gonna cancel

OpenStudy (dumbcow):

First, factor 9a^2 -b^2 (3a-b)(3a+b) Then Flip and Multiply when dividing fractions flip the bottom fraction and change to multiplication \[\frac{(3a-b)(3a+b)}{ab}*\frac{a}{3a+b}\] Now cancel the a on top and bottom cancel the (3a+b) on top and bottom leaving.. (3a-b)/b

OpenStudy (anonymous):

wow, the cow did it again i am impressed

OpenStudy (anonymous):

medal for you!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

it was hard to see at first, but I understand the question. Cool beans.

OpenStudy (dumbcow):

and to you for helping

OpenStudy (anonymous):

well its late. The monkey needs its rest. Goodnight everyone.

OpenStudy (dumbcow):

tina does it make sense? ... goodnight

OpenStudy (gina):

(3a+b)(3a-b)=9a^2-b^2